### v2

# Learning classical density functionals for ionic fluids

Abstract
Accurate and efficient theoretical techniques for describing ionic fluids are highly desirable for many applications across the physical, biological and materials sciences. With a rigorous statistical mechanical foundation, classical density functional theory (cDFT) is an appealing approach, but the competition between strong Coulombic interactions and steric repulsion limits the accuracy of current approximate functionals. Here, we extend a recently presented machine learning (ML) approach [Sammüller et al., Proc. Natl. Acad. Sci. USA, 120, e2312484120 (2023)] designed for systems with short-ranged interactions to ionic fluids. By adopting ideas from local molecular field theory, the framework we present amounts to using neural networks to learn the local relationship between the one-body direct correlation functions and inhomogeneous density profiles for a “mimic” short-ranged system, with effects of long-ranged interactions accounted for in a mean-field, yet well-controlled, manner. By comparing to results from molecular simulations, we show that our approach accurately describes the structure and thermodynamics of prototypical models for electrolyte solutions and ionic liquids, including size-asymmetric and multivalent systems. The framework we present acts as an important step toward extending ML approaches for cDFT to systems with accurate interatomic potentials.
The behavior of ionic fluids underlies a vast array of physical and biological phenomena as well as technological applications, ranging from electrolyte solutions controlling protein folding [1] to room-temperature ionic liquids in energy storage devices [2]. A fundamental topic that continues to attract enormous interest both experimentally [3, 4, 5, 6, 7] and theoretically [8, 9, 10, 11, 12, 13] is the structure and thermodynamics of ions near charged interfaces, in particular, how the nature of both the electrolyte and solid surface impacts the properties of the electric double layer (EDL). Poisson–Boltzmann (PB) theory and its linearized Debye-Hückel form provide the basis for much of our understanding of ionic fluids. Their neglect of correlations arising from non-electrostatic interactions, however, restricts their validity to fluids of low ionic strength.
Classical density functional theory (cDFT) [14, 15, 16, 17] provides a natural framework for including correlations omitted by PB theory, and has proven to be a powerful approach to describe equilibrium structure and thermodynamics of fluids in general. While in principle an exact theory, historically, cDFT relies on making approximations for the excess intrinsic free energy functional $\mathcal{F}^{\mathrm{(ex)}}_{\mathrm{intr}}[\{\rho_{\nu}\}]$, where $\rho_{\nu}$ denotes the one-body density of species $\nu$. For hard sphere fluids, functionals based on Rosenfeld’s fundamental measure theory (FMT) [18, 19, 20, 21] have proven highly successful. For simple liquids with square-well, Yukawa, or Lennard–Jones interaction potentials, hard sphere mixtures act as suitable reference systems, with effects of attractive interactions described reasonably well in a mean-field fashion [22, 23, 24]. In contrast, existing functionals for ionic fluids are far less accurate, failing to adequately capture the interplay between Coulombic and steric interactions [25, 26, 27, 28]. In this Letter, we present a strategy that utilizes machine learning (ML) to construct accurate free energy functionals for ionic fluids.
Figure 1: Structure of the SR mimic RPM. (a) Training data for the neural functional are obtained from GCMC simulations with (i) insertion/deletion, (ii) position swapping, (iii) displacement and (iv) mutation (identity exchange) moves. (b) The static structure factors from cDFT agree well with results from molecular simulation across the whole range of $k$ for the SR system, shown for a bulk system with $\sigma^{3}\rho_{\mathrm{R},\pm}=0.315$. Inset: At low $k$, the SR system violates the perfect screening condition, whereas the LR system obeys the Stillinger–Lovett sum rule (see Eq. 5). (c) For the applied external potentials (top), predictions of the ion density profiles (bottom) are in excellent agreement with the simulation data.
The rapid advance of modern ML approaches means there has been much recent interest in “learning” representations of the exact $\mathcal{F}^{\mathrm{(ex)}}_{\mathrm{intr}}[\{\rho_{\nu}\}]$ [29, 30, 31, 32, 33, 34, 35]. Here we build on the method proposed by Sammüller et al. [36], in which the one-body direct correlation functions,
$c^{(1)}_{\nu}(\bm{\mathit{r}};[\{\rho_{\nu}\}])=-\frac{\beta\delta\mathcal{F}^{\mathrm{(ex)}}_{\mathrm{intr}}[\{\rho_{\nu}\}]}{\delta\rho_{\nu}(\bm{\mathit{r}})},$ (1)
are learned by generating inhomogeneous density profiles in the presence of various external- and chemical potentials by grand canonical (GC) simulations. This approach to cDFT, dubbed “neural functional theory,” has been shown to outperform FMT-based approaches for both hard sphere fluids and Lennard–Jones liquids in terms of both accuracy and speed [36, 37].
Two key features underpin the success of the neural functional approach: (i) correlations in the fluid are short-ranged (SR), such that the functional relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ is local; and (ii) the feasibility of GC simulations to produce inhomogeneous density profiles of sufficient quality to form a reliable training set. Ionic fluids pose challenges on both fronts. First, the long-ranged (LR) nature of the Coulomb potential leads to a nonlocal relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$. Second, GC simulations for ionic systems raise delicate questions concerning electroneutrality, with various schemes proposed that either insert individual ions or neutral pairs [38, 39, 40, 41, 42]. Even then, simulations corresponding exactly to the GC ensemble where the number of each species can fluctuate are impractical, owing to poor acceptance rates of trial insertion and deletion moves. We circumvent both of these issues by adopting concepts from local molecular field theory (LMFT) [43, 44, 45, 46, 47], which has been shown to have close links to cDFT [48].
We initially focus our efforts on the prototypical model of an ionic fluid: the restricted primitive model (RPM) comprising oppositely charged hard spheres of equal diameter $\sigma$, embedded in a uniform dielectric continuum. Later, we will also present results for a primitive model (i.e., size-asymmetric) and a multivalent system. The scheme we propose can be briefly summarized. First, by employing neural networks, we find the one-body direct correlation functions for a suitably chosen “mimic system” whose electrostatic interactions are entirely short-ranged. Then, by leveraging LMFT and its relation to cDFT, we account for the net averaged effects of LR electrostatic interactions in a well-controlled fashion. When compared to molecular simulations, the framework we outline describes inhomogeneous density profiles, the equation of state, and the properties of the electric double layer in the presense of electric fields with very high accuracy.
cDFT of a short-ranged “mimic” ionic fluid. Our overall strategy follows that of LMFT, in which we consider a suitably chosen mimic system whose interatomic interactions are entirely short-ranged, and, when subject to a suitably chosen one-body potential $\phi_{\rm R}$, has the same one-body densities as the system of interest with LR interactions (the “full” system). For systems such as the RPM, where LR interactions arise from the Coulomb potential, one adopts the exact splitting
$1/r=v_{0}(r)+v_{1}(r),$ (2)
with $v_{0}(r)=\mathrm{erfc}(\kappa r)/r$ and $v_{1}(r)=\mathrm{erf}(\kappa r)/r$. The interatomic potential of the mimic system is then $v_{0}$, and the length scale $\kappa^{-1}$ is chosen such that the mimic system accurately describes the SR correlations of the full system. In the following, we will work with $\kappa^{-1}=1.8\,\sigma$ (see the SM [49]). We will also indicate quantities pertaining to the mimic system with a subscript “R” and focus on behaviors at a reduced temperature $T^{*}=0.066$, corresponding to supercritical conditions  [^1] The critical temperature of the RPM is $T^{*}_{\mathrm{c}}=0.0492$ (from Ref. 39). The one-body direct correlation functions of the mimic system can be exactly written as
$c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}})=\ln\Lambda_{\nu}^{3}\rho_{\mathrm{R},\nu}(\bm{\mathit{r}})+\beta V_{\mathrm{R},\nu}(\bm{\mathit{r}})+\beta q_{\nu}\phi_{\mathrm{R}}(\bm{\mathit{r}})-\beta\mu_{\mathrm{R},\nu},$ (3)
where $\Lambda_{\nu}$, $\mu_{\mathrm{R},\nu}$ and $q_{\nu}$ indicate the thermal de Broglie wavelength, chemical potential and point charge of species $\nu$, respectively, $V_{\mathrm{R},\nu}$ encompasses any non-electrostatic contributions to the external potential for species $\nu$ and, for now, $\phi_{\mathrm{R}}$ is a general external electrostatic potential; we will later discuss how $\phi_{\rm R}$ can be chosen such that $\rho_{\mathrm{R},\nu}(\bm{\mathit{r}})=\rho_{\nu}(\bm{\mathit{r}})$.
To learn the functional relationship for $c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}};[\{\rho_{\mathrm{R},\nu}\}])$, we obtain data for the right hand side of Eq. 3 by measuring $\rho_{\mathrm{R},\nu}$ from simulations with known $\beta V_{\mathrm{R},\nu}$, $\beta\phi_{\mathrm{R}}$ and $\beta\mu_{\mathrm{R},\nu}$. To this end, in line with Ref. 36, we perform GC simulations in a planar geometry at different combinations of $\{\beta V_{\mathrm{R},\nu}\}$, $\beta\phi_{R}$ and $\{\beta\mu_{\mathrm{R},\nu}\}$. Note that GC simulations are essential for the purpose of evaluating Eq. 3 and, on their own, canonical methods such as molecular dynamics (MD) simulations are insufficient. The form of $v_{0}$ means that each particle of the mimic system can be considered electroneutral, comprising both a point charge and a compensating Gaussian charge distribution [51]. As such, in addition to translational moves, we can readily perform GC particle insertions/deletions, along with semi-GC swapping of “anions” and “cations” without needing to worry about issues of electroneutrality [see Fig. 1(a)]. We have found this highly beneficial to converging our simulations, full details of which are provided in the SM [49]. In total, $\sim 2500$ simulations have been performed to gather training data.
For each species, $\nu=+\,\mathrm{or}\,-$, the neural network used to represent the relationship $\{\rho_{\mathrm{R},\nu}(z)\}\rightarrow c^{(1)}_{\mathrm{R},\nu}(z)$ is structured as follows. The input layer has two channels that are supplied with the discretized values of the cation and anion density profiles in a window $\Delta z=3.6\,\sigma$ around the particular value of $z$. Following a fully-connected multilayer perceptron of three layers, the output layer consists of a single node, which yields the predicted value of $c^{(1)}_{\mathrm{R},\nu}$ at position $z$. For each ionic fluid considered, we train two independent networks, one for the cation and one for the anion. To reduce noise in the bulk structure predictions, we also obtained models regularized with bulk two-body direct correlation functions as proposed in Ref. 52. Details of our training procedure are provided in the SM [49], and all data are openly available [53].
Figure 2: Structure and thermodynamics of the LR full system. cDFT for the RPM at $T^{*}=0.066$ with effects of LR electrostatic interactions accounted by LMFT shows excellent agreement with reference molecular simulation data for (a) the equation of state, (b) ion density profiles confined in a slit and (c) with an applied external potential. In (a), we see that our cDFT approach outperforms the prediction obtained by combining the Carnahan–Starling equation of state for hard-sphere fluids ($P^{\mathrm{CS}}$) and the mean spherical approximation ($P^{\mathrm{MSA}}$). For the LR system in (c), the applied external potential acts over a much longer range and the chemical potentials are shifted lower compared to the SR mimic system in Fig. 1(c) to yield $\rho_{\nu}(z)=\rho_{\mathrm{R},\nu}(z)$.
We first assess the bulk structure predicted by the functional. On the basis of the neural network, we make use of automatic differentiation and radial projection to obtain the partial two-body direct correlation functions $c^{(2)}_{\nu\lambda}(r)$ (see Ref. [36] and the SM [49]). By solving the general Ornstein–Zernike equation for mixtures [54, 55], we then obtain the total correlation functions $h^{(2)}_{\nu\lambda}(r)$ and the corresponding structure factors $S_{\nu\lambda}(k)$. We further quantify the degree of coupling between number–number ($S_{NN}$), number–charge ($S_{NZ}$) and charge–charge ($S_{ZZ}$) densities by appropriate weighted summations
$\begin{split}S_{NN}(k)&=S_{++}(k)+2S_{+-}(k)+S_{--}(k),\\ S_{NZ}(k)&=S_{++}(k)-S_{--}(k),\\ S_{ZZ}(k)&=S_{++}(k)-2S_{+-}(k)+S_{--}(k).\end{split}$ (4)
Fig. 1(b) shows that the static structure factors obtained from the functional are in excellent agreement with results from a molecular dynamics simulation of the mimic system at the same bulk densities  [^2] While GC simulations (i.e., known $\{\mu_{\nu}\}$) are required to train the neural networks (see Eq. 3), it is reasonable to compare to average structures obtained with canonical simulations at the same average density.. The symmetry of the RPM is well-captured, reflected in $S_{NZ}(k)=0$. Moreover, by comparing to results from a MD simulation of the full system, we see that $S_{NN}(k)$ and $S_{ZZ}(k)$ for the full and mimic systems agree very well at sufficiently large $k$, confirming that our choice of $\kappa^{-1}=1.8\,\sigma$ is sufficient for the mimic system to capture the SR correlations of the full system. Deviations of the SR system from the LR system only manifest significantly in $S_{ZZ}(k)$ at small $k$, in agreement with previous works on the subject [46, 57]. In particular, as shown in the inset, we see that $S_{ZZ}(k)$ for the LR system strictly obeys the Stillinger–Lovett sum rule [58, 59]
$\lim_{k\rightarrow 0}\frac{k^{2}_{\rm D}}{k^{2}}S_{ZZ}(k)=1,$ (5)
where $k_{\rm D}^{-1}$ is the Debye screening length. In contrast, for the mimic system we observe $S_{ZZ}(k)>k^{2}/k^{2}_{\rm D}$ as $k\to 0$, indicative of a lack of screening. Consistent with the dielectric response of a short-ranged water model [57], these deviations appear at length scales far larger than the range separation prescribed by $2\pi\kappa=3.5\,\sigma^{-1}$.
Turning to the ability of cDFT to predict inhomogeneous structure, the equilibrium density profiles of the mimic system can be obtained by self-consistently solving the Euler–Lagrange equation
$\begin{split}\Lambda^{3}_{\nu}\rho_{\mathrm{R},\nu}(\bm{\mathit{r}})=\exp\Big{(}-\beta V_{\mathrm{R},\nu}(\bm{\mathit{r}})&-\beta q_{\nu}\phi_{R}(\bm{\mathit{r}})\\ &+\beta\mu_{\mathrm{R},\nu}+c^{(1)}_{\mathrm{R},\nu}(\bm{\mathit{r}};[\{\rho_{\mathrm{R},\nu}\}])\Big{)},\end{split}$ (6)
where $c_{\mathrm{R},\nu}^{(1)}$ is evaluated by the corresponding neural network. As shown in Fig. 1(c) for a representative set of external potentials, the resulting $\{\rho_{\mathrm{R},\nu}(z)\}$ from cDFT are in excellent agreement with reference simulation data of the mimic system.
Accounting for long-ranged electrostatics with LMFT. Having established that the cDFT obtained from the ML procedure is accurate for a SR variant of the RPM, we turn our attention to incorporating the effects of LR electrostatic interactions. To do so, we will use concepts from LMFT. The premise of LMFT is that there exists a potential $\phi_{\rm R}$ such that
$\rho_{\nu}(\bm{\mathit{r}};[\phi],\{\mu_{\nu}\})=\rho_{\mathrm{R},\nu}(\bm{\mathit{r}};[\phi_{\mathrm{R}}],\{\mu_{\mathrm{R},\nu}\}).$ (7)
In Eq. 7, we have explicitly indicated the functional dependence of the densities on the external electrostatic potential  [^3] The number densities also have a functional dependence on the non-electrostatic contribution to the external potential. As this is the same for both the full and mimic systems, i.e $\{V_{\nu}\}=\{V_{\nu,\mathrm{R}}\}$, we do not indicate this functional dependence explicitly in Eq. 7.. As shown in Ref. 48, when recast in a cDFT framework, LMFT relates the one-body direct correlation functions of the full and mimic systems through
$c_{\nu}^{(1)}(\bm{\mathit{r}};[\{\rho_{\nu}\}])=c_{\mathrm{R},\nu}^{(1)}(\bm{\mathit{r}};[\{\rho_{\nu}\}])-\beta\Delta\mu_{\nu}+\beta q_{\nu}\Delta\phi(\bm{\mathit{r}}),$ (8)
which is an exact result. A key insight from LMFT [44, 43] is that for an appropriate splitting of the potential (see Eq. 2), $\Delta\phi$ is well-approximated by a mean field form
$\Delta\phi(\bm{\mathit{r}})\equiv\phi(\bm{\mathit{r}})-\phi_{\mathrm{R}}(\bm{\mathit{r}})=-\frac{1}{\epsilon}\int\!\!\mathrm{d}\bm{\mathit{r}}^{\prime}\,n(\bm{\mathit{r}}^{\prime})\,v_{1}(|\bm{\mathit{r}}-\bm{\mathit{r}}^{\prime}|),$ (9)
where $\epsilon$ is the dielectric constant of the continuum and $n=n_{\mathrm{R}}=\sum_{\nu}q_{\nu}\rho_{\nu}$ is the charge density. For homogeneous systems, Rodgers and Weeks have also derived a thermodynamic correction for the total energy, from which a correction for the pressure follows [61],
$\Delta P\equiv P-P_{R}=-\frac{k_{B}T}{2\pi^{3/2}\kappa^{-3}}.$ (10)
Based on Ref. 61, we have further derived a correction for the chemical potential (see the SM [49]),
$\Delta\mu_{\nu}\equiv\mu_{\nu}-\mu_{\mathrm{R},\nu}=-\frac{q^{2}_{\nu}}{\kappa^{-1}\sqrt{\pi}}.$ (11)
We first consider the bulk thermodynamics of the RPM. The equation of state for the LR system is obtained with
$P(\{\rho_{\nu}\})=\sum_{\nu}k_{\mathrm{B}}T\rho_{\nu}\left(1-c^{(1)}_{\mathrm{R},\nu}[\{\rho_{\nu}\}]\right)-\frac{\mathcal{F}^{\mathrm{(ex)}}_{\mathrm{intr},\mathrm{R}}[\{\rho_{\nu}\}]}{V}+\Delta P,$ (12)
where the sum of the first two terms is the negative of the grand potential density of the mimic system. The excess free energy $\mathcal{F}^{\mathrm{(ex)}}_{\mathrm{intr},\mathrm{R}}$ is accessible by functional line integration. The result of this procedure, shown in Fig. 2(a), agrees very well with reference simulation data. We also see that, particularly at higher densities, Eq. 12 performs significantly better than the analytical approximation that results from adding the Carnahan–Starling equation of state ($P^{\mathrm{CS}}$) [62] and the mean spherical approximation ($P^{\mathrm{MSA}}$) [63, 64], despite the fact that $P^{\mathrm{CS}}+P^{\mathrm{MSA}}$ forms the basis for the vast majority of current state-of-the-art functionals for ionic fluids [27, 28].
The advantages of cDFT combined with LMFT become even clearer when inhomogeneous systems are considered. The equilibrium densities of the full system are given by
$\begin{split}\Lambda^{3}_{\nu}\rho_{\nu}(\bm{\mathit{r}})=\exp\Big{(}-\beta V_{\nu}(\bm{\mathit{r}})&-\beta q_{\nu}\phi(\bm{\mathit{r}})\\ &+\beta\mu_{\nu}+c^{(1)}_{\nu}(\bm{\mathit{r}};[\{\rho_{\nu}\}])\Big{)},\end{split}$ (13)
where the non-local functional $c_{\nu}^{(1)}$ is now given by Eq. 8, together with Eq. 9 and Eq. 11. From a cDFT perspective, we have captured all strong rapidly-varying short-ranged correlations with the neural networks for $c_{\mathrm{R},\nu}^{(1)}$, while the effects of LR electrostatic interactions are incorporated in a mean-field, yet well-controlled fashion.
In practical terms, for a planar geometry, Eq. 9 can be recast as
$\Delta\phi(z)=-\frac{1}{\epsilon L}\sum_{k\neq 0}\frac{4\pi}{k^{2}}\tilde{n}(k)\exp(ikz)\exp\left(-\frac{k^{2}}{4\kappa^{2}}\right),$ (14)
where $\tilde{n}$ denotes a Fourier component of $n$, and $L$ is the total length of the periodic cell. In Fig. 2(b), we show the results from our cDFT approach for the RPM confined between two repulsive walls for various values of $\mu_{+}=\mu_{-}$. The density profiles are in very good agreement with reference canonical MD simulation data in which the number of particles matches that obtained by integrating the density profiles from cDFT. Our cDFT approach also works very well in cases where the external potential contains an electrostatic component, as illustrated in Fig. 2(c). Here, the full system is the LR counterpart of the mimic system shown in Fig. 1(c), emphasizing the premise of LMFT (see Eq. 7).
The electric double layer and other ionic fluids. The results presented so far demonstrate that our cDFT approach provides an efficient and accurate route to the structure and thermodynamics of the RPM. As an application, we turn our attention to the fundamental topic of ongoing scientific interest: the electric double layer. We also show that our approach is robust to choice of interatomic potential. Specifically, we also consider the primitive model (PM) and a multivalent fluid.
As a simple model of an EDL-forming system, we consider the RPM confined between two repulsive walls, and in the presence of an electric field $E_{z}$ along the $z$ direction. The electrostatic potential that the LR system feels is $\phi(z)=-E_{z}z$. For the reference simulation data, the LR system was simulated at constant $E_{z}$ using the finite-field approach [65, 66, 67, 68, 69, 70, 71]. As shown in Fig. 3(a), the ion distributions from our cDFT approach are in excellent agreement with the results from simulation. Notably, the cDFT accurately describes the strong oscillations in the density profiles of co- and counter-ions. Moreover, as we detail in the SM [49], our functional is thermodynamically consistent, satisfying the contact density theorem [72, 73, 74]
$P=-\sum_{\nu}\int\!\!\mathrm{d}z\,\rho_{\nu}(z)\frac{\mathrm{d}V_{\nu}(z)}{\mathrm{d}z}-\frac{2\pi\sigma_{\mathrm{s}}^{2}}{\epsilon},$ (15)
which relates the bulk pressure to the surface charge density $\sigma_{\rm s}=\int_{-L}^{0}\!\mathrm{d}z\,n(z)$  [^4] This equation for $\sigma_{\rm s}$ assumes that the electrolyte is centered in a periodically replicated cell of length $L$ along the $z$ direction. and ion densities at contact with the walls. Obeying this contact theorem has proven a challenge not only for integral equation methods [76, 77, 78, 79], which tend to violate key sum rules and are therefore generally thermodynamically inconsistent [80, 48], but also for most state-of-the-art cDFT functionals for ionic fluids [81, 82]. The accurate description of EDL structure and thermodynamic consistency displayed by our LMFT-based neural cDFT approach constitutes a significant advancement in the theoretical description of ionic fluids.
Figure 3: Accurately describing the EDL with cDFT for (a) the RPM, (b) the PM, and (c) a multivalent fluid. In all cases, confining walls located at ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z/\sigma=\pm 7.2}$ in a periodic cell of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L/\sigma=18.1}$, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{z}=1.7\,k_{\mathrm{B}}T/e\sigma}$. The resulting ion density profiles from cDFT are in excellent agreement with molecular simulations, and obey the contact theorem (Eq. 15).
The main advantage of the neural functional approach outlined in Ref. 36 is arguably that the local relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ permits application of the resulting functional to system sizes far beyond those encountered during training of the ML. By using the framework of LMFT and its relationship to cDFT, we have successfully extended the neural functional approach to a case where the relationship between $c_{\nu}^{(1)}$ and $\{\rho_{\nu}\}$ is nonlocal. To demonstrate that this methodology is not limited to the RPM, in Figs. 3(b) and (c), we present results for a PM and multivalent system. For the PM, we have changed the sizes of the anion and cation to $\sigma_{-}=4\sigma/3$ and $\sigma_{+}=2\sigma/3$, respectively. For the multivalent system, we consider equal-sized anions and cations of diameter $\sigma$, but increase the valency of the anion such that $q_{-}/q_{+}=2$. We use the same temperature as for the RPM, $T^{*}=\epsilon\sigma k_{\mathrm{B}}T/|q_{+}q_{-}|=0.066$. In both cases, we observe excellent agreement between the neural cDFT and simulations, both for inhomogeneous profiles and bulk equations of state, as further shown in the SM [49]). Future work will pursue using neural functionals to understand EDL capacitance [83] and surface force balance measurements [13], and to augment other cDFT approaches to understand solvation [84]. The framework we have outlined should also find use in describing the dielectric response of polar fluids such as water [85, 86, 57].
Data availability. Training data and code supporting the findings of this study will be openly available [103, 53].
[…]